Starting from rest, a car accelerates at 2.2 m/s2 up a hill that is inclined 5.5° above the horizontal. How far horizontally and vertically has the car traveled in 12 s?
Answer #1
Kinematics equation S = ut +1/2 at^2
u = initial velocity = 0 m/s
Thus, S = 1/2at^2
Substituting the numerical values,
S = (0.5)(2.2 m/s^2)(12s)^2 =158.4 m
Now let's find how far the car has traveled horizontally and vertically
Horizontally it would have traveled =(158.4 m)(cos 5.5 degrees) = 157.6 m
Vertically it would have traveled =(158.4 m)(sin 5.5 degrees) = 15.2 m
Answer #2

Break the acceleration of into horizontal and vertical components.

uploaded imagenext step

The car starts from rest.